# Bilevel Reinforcement Learning via the Development of Hyper-gradient without Lower-Level Convexity

### v1

@@ G.4 Convergence properties of $\pi_{k}$ and $Q_{k}$
The softmax mapping, plays a significant role in the update rule of $\pi_{k}$ and $Q_{k}$, for which we begin this section by introducing some properties of it, following from [9]. Given a $\theta\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$, for any $s\in\mathcal{S}$, use $\theta_{s}\in\mathbb{R}^{\left|{\mathcal{A}}\right|}$ to denote a vector with $\theta_{s}(a)=\theta_{sa}=\theta(s,a)$. Recall the softmax mapping:
@@
We present the properties of the operator $\mathcal{T}_{\mathcal{M}_{\tau}\left(x\right)}:\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}\mapsto\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$, following from [29, 9].
@@
Substitute one of the consistency conditions from [62],
@@ G.5 Convergence analysis of M-SoBiRL
Under Assumptions 6.1 and 6.3, the following inequalities hold in Algorithm 1.
@@
$\leq$ $\ \sqrt{\left|\mathcal{S}\right|}C_{rx}\left\|w_{k}-w_{k}^{*}\right\|_{2}+\frac{2\left|\mathcal{S}\right|\left|\mathcal{A}\right|^{3/2}\sqrt{\left(1+\gamma\right)}C_{f\pi}C_{rx}}{\tau\left(1-\gamma\right)}\left\|V_{k}^{*}-V_{k}\right\|_{\infty}$
@@
Under Assumptions 6.1 and 6.3, the hyper-gradient estimator constructed in Algorithm 1,
@@
    $\delta_{w}^{k+1}\leq$ $\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}+4\xi^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\right)\delta_{w}^{k}+\frac{4\xi^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\delta_{Q}^{k+1}$
    $+2L_{w}^{2}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)\beta^{2}\left\|\widehat{\nabla}\phi_{k}\right\|_{2}^{2}$ (48)
@@
Under Assumptions 6.1 and 6.3, in Algorithm 1, we can choose constant step sizes $\beta,\xi$, and the inner iteration number $N\sim\mathcal{O}(1)$. Then the iterates $\{x_{k}\}$ satisfy
@@
$+2\gamma^{2N}\left(\zeta_{Q}+\frac{4\xi^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\zeta_{w}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)\left(\delta_{Q}^{k}+\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\beta^{2}\left\|\widehat{\nabla}\phi_{k}\right\|_{2}^{2}\right)$
$+\zeta_{w}\left(1-\frac{\xi\left(1-\gamma\right)^{2}}{2}+4\xi^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\right)\delta_{w}^{k}$
$+2L_{w}^{2}\zeta_{w}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)\beta^{2}\left\|\widehat{\nabla}\phi_{k}\right\|_{2}^{2}-\zeta_{Q}\delta_{Q}^{k}-\zeta_{w}\delta_{w}^{k}$ (53)
@@
$I_{w}=$ $\ \frac{3\beta\left|\mathcal{S}\right|C_{rx}^{2}}{2}+\zeta_{w}\left(-\frac{\xi\left(1-\gamma\right)^{2}}{2}+4\xi^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\right)$
$I_{Q}=$ $\ 2\gamma^{2N}\left(\zeta_{Q}+\frac{4\xi^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\zeta_{w}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)-\zeta_{Q}$
@@
$I_{\widehat{\phi}}=$ $\ \frac{L_{\phi}^{M}}{2}\beta^{2}+2\beta^{2}L_{w}^{2}\zeta_{w}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)-\frac{\beta}{2}$
$+2\gamma^{2N}\left(\zeta_{Q}+\frac{4\xi^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\zeta_{w}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\beta^{2},$
@@
$\begin{cases}\zeta_{Q}=\zeta_{w}=1,\\ \frac{3\beta\left|\mathcal{S}\right|C_{rx}^{2}}{2\xi\left(1-\gamma\right)^{2}}<\frac{1}{2}-4\xi\left|\mathcal{S}\right|^{2}\frac{\left(1+\gamma\right)^{2}}{\left(1-\gamma\right)^{2}},\\ \beta<\frac{\tau^{2}}{8}\left(L_{\widehat{\phi}\pi}^{2}+\frac{\left|\mathcal{S}\right|^{2}\left|\mathcal{A}\right|^{3}\left(1+\gamma\right)C_{f\pi}^{2}C_{rx}^{2}}{\left(1-\gamma\right)^{2}}\right)^{-1},\\ \gamma^{2N}<\frac{1}{8}\left(1+\frac{4\xi^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)^{-1},\\ \frac{L_{\phi}^{M}}{2}\beta+2\beta L_{w}^{2}\left(1+\frac{1}{\xi\left(1-\gamma\right)^{2}}\right)+\frac{1}{4}\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\beta<\frac{1}{2}.\end{cases}$ (55)
Set the step size $\beta=\rho\xi$ with $\rho>0$. More precisely, we present the parameter configuration to guarantee (55).
$\begin{cases}\zeta_{Q}=\zeta_{w}=1,\\ \xi<\min\left\{1,\frac{\left(1-\gamma\right)^{2}}{16\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}}\right\},\\ \rho<\min\left\{\frac{\left(1-\gamma\right)^{2}}{6\left|\mathcal{S}\right|C_{rx}^{2}},\frac{\left(1-\gamma\right)^{2}}{8L_{w}^{2}}\right\},\\ \beta<\min\left\{\frac{\tau^{2}}{8}\left(L_{\widehat{\phi}\pi}^{2}+\frac{\left|\mathcal{S}\right|^{2}\left|\mathcal{A}\right|^{3}\left(1+\gamma\right)C_{f\pi}^{2}C_{rx}^{2}}{\left(1-\gamma\right)^{2}}\right)^{-1},\frac{1}{4}\left(\frac{L_{\phi}^{M}}{2}+2L_{w}^{2}+\frac{1}{4}\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\right)^{-1}\right\},\\ \gamma^{2N}<\frac{1}{8}\left(1+\frac{4C_{\sigma\pi}^{2}}{\tau^{2}}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)^{-1},\end{cases}$ (56)
@@ Appendix H Analysis of SoBiRL
The next proposition, generalized from Lemma 1 in [10], considers the distribution $\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi\right)$, with each trajectory $d_{i}=\{\left({s_{h}^{i},a_{h}^{i}}\right)\}_{h=0}^{H-1}$ ($i=1,2,\ldots,I$) sampled from the trajectory distribution $\rho\left({d;\pi}\right)$, i.e.,
@@
Denote the total variation between distributions by $D_{F}\left({\cdot,\cdot}\right)$. For any trajectory tuple $\left(d_{1},d_{2},...,d_{I}\right)$, with each trajectory holding a finite horizon $H$, we have
$D_{F}\left(\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi^{1}\right),\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi^{2}\right)\right)\leq\frac{HI}{2}\sqrt{\left|\mathcal{A}\right|}\left\|\pi^{1}-\pi^{2}\right\|_{2}.$ (57)
@@
$D_{F}\left(\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi^{*}(x_{1})\right),\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi^{*}(x_{2})\right)\right)\leq\frac{HIL_{\pi}}{2}\sqrt{\left|\mathcal{A}\right|}\left\|x_{1}-x_{2}\right\|_{2}.$
@@
Note that $\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi\right)$ is the product measure of $\rho\left(d_{i};\pi\right),\,\left(i=1,2,...,I\right)$, i.e., $\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi\right)=\rho\left(d_{1};\pi\right)\times\rho\left(d_{2};\pi\right)\times\cdots\times\rho\left(d_{I};\pi\right)$. By the total variation inequality for the product measure,
$D_{F}\left(\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi^{1}\right),\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi^{2}\right)\right)\leq$ $\sum_{i=1}^{I}{D_{F}\left(\rho\left(d_{i};\pi^{1}\right),\rho\left(d_{i};\pi^{2}\right)\right)}$
@@
If the vector-valued function $z\left({d_{1},d_{2},\ldots,d_{I};x}\right)\in\mathbb{R}^{n}$ is bounded by $C$, i.e., $\left\|{z\left({d_{1},d_{2},\ldots,d_{I};x}\right)}\right\|_{2}\leq C$, then for any trajectory tuple $\left(d_{1},d_{2},...,d_{I}\right)$, with each trajectory holding a finite horizon $H$, and policies $\pi_{1},\pi_{2}$, it holds
@@
We resort to the property of the Integral Probability Metric (IPM) [73],
@@
$\left|b^{\top}\left(\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{1}\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{2}\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]\right)\right|$
$\leq$ $\ 2C\left\|b\right\|_{2}D_{F}\left(\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi_{1}\right),\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi_{2}\right)\right)$
@@
$\left\|{\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{1}\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{2}\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]}\right\|_{2}$
$\leq$ $\ 2C\times D_{F}\left(\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi_{1}\right),\rho\left(\left(d_{1},d_{2},...,d_{I}\right);\pi_{2}\right)\right).$ (64)
@@
$Z\left(x_{1}\right)-Z\left(x_{2}\right)=$ $\left(\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{1}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]\right)$
$+\left(\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{2}\right)\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]\right).$
@@
$\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{2}\right)\right)}\left[z\left(d_{1},d_{2},...,d_{I};x_{2}\right)\right]\right\|\leq CHIL_{\pi}\sqrt{\left|{\mathcal{A}}\right|}\left\|x_{1}-x_{2}\right\|_{2}.$
@@
The following proposition measures the quality of the implicit differentiation estimators. Similar result is introduced as an assumption in [69] (Lemma 18.condition (c)). We claim that it is satisfied in our setting.
@@
Under Assumption 6.3, for any upper-level variable $x\in\mathbb{R}^{n}$ and policies $\pi^{1},\pi^{2}$, we have for any $s\in\mathcal{S}$, $a\in\mathcal{A}$,
$\max_{s,a}\left\|\widetilde{\nabla}Q_{sa}\left({x,\pi^{1}}\right)-\widetilde{\nabla}Q_{sa}\left(x,\pi^{2}\right)\right\|_{2}$ $\leq\frac{(1+\gamma)C_{rx}\sqrt{\left|\mathcal{A}\right|}}{\left(1-\gamma\right)^{2}}\left\|\pi^{1}-\pi^{2}\right\|_{2}$
@@
Under the Assumptions 6.2 and 6.3, $\nabla\phi(x)$ is $L_{\phi}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$,
@@
$\nabla\phi(x)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},...,d_{I};x\right)\right]$ (69)
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right],$
@@
$z\left({d_{1},d_{2},\ldots,d_{I};x}\right):=l\left(d_{1},d_{2},...,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)$
@@
$\left\|z\left(d_{1},d_{2},...,d_{I};x\right)\right\|_{2}\leq 2HIC_{l}\frac{C_{rx}}{1-\gamma},$
@@
Under Assumptions 6.2 and 6.3, for any upper-level variable $x\in\mathbb{R}^{n}$ and policies $\pi^{1},\pi^{2}$, we have

### v2

@@ G.4 Convergence Properties of $\pi_{k}$ and $Q_{k}$
The softmax mapping, plays a significant role in the update rule of $\pi_{k}$ and $Q_{k}$, for which we begin this section by introducing some properties of it, following from (Cen et al.,, 2022). Given a $\theta\in\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$, for any $s\in\mathcal{S}$, use $\theta_{s}\in\mathbb{R}^{\left|{\mathcal{A}}\right|}$ to denote a vector with $\theta_{s}(a)=\theta_{sa}=\theta(s,a)$. Recall the softmax mapping:
@@
We present the properties of the operator $\mathcal{T}_{\mathcal{M}_{\tau}\left(x\right)}:\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}\mapsto\mathbb{R}^{\left|{\mathcal{S}}\right|\left|{\mathcal{A}}\right|}$, following from (Haarnoja et al., 2018a, ; Cen et al.,, 2022).
@@
Substitute one of the consistency conditions from (Nachum et al.,, 2017),
@@ G.5 Convergence Analysis of M-SoBiRL
Under Assumptions 7.1 and 7.3, the following inequalities hold in Algorithm 2.
@@
$\leq$ $\ \sqrt{\left|\mathcal{S}\right|}C_{rx}\left\|w_{k}-w_{k}^{*}\right\|_{2}+\frac{2\left|\mathcal{S}\right|\left|\mathcal{A}\right|^{3/2}\sqrt{\left(1+\gamma\right)}C_{f\pi}C_{rx}}{\tau\left(1-\gamma\right)}\left\|V_{k}^{*}-V_{k}\right\|_{\infty}.$
@@
Under Assumptions 7.1 and 7.3, the hyper-gradient estimator constructed in Algorithm 2,
@@
    $\delta_{w}^{k+1}\leq$ $\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}+4\eta^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\right)\delta_{w}^{k}+\frac{4\eta^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\delta_{Q}^{k+1}$
    $+2L_{w}^{2}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)\beta^{2}\left\|\widehat{\nabla}\phi_{k}\right\|_{2}^{2}$ (48)
@@
Under Assumptions 7.1 and 7.3, in Algorithm 2, we can choose constant step sizes $\beta,\eta$, and the inner iteration number $N\sim\mathcal{O}(1)$. Then the iterates $\{x_{k}\}$ satisfy
@@
$+2\gamma^{2N}\left(\zeta_{Q}+\frac{4\eta^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\zeta_{w}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)\left(\delta_{Q}^{k}+\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\beta^{2}\left\|\widehat{\nabla}\phi_{k}\right\|_{2}^{2}\right)$
$+\zeta_{w}\left(1-\frac{\eta\left(1-\gamma\right)^{2}}{2}+4\eta^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\right)\delta_{w}^{k}$
$+2L_{w}^{2}\zeta_{w}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)\beta^{2}\left\|\widehat{\nabla}\phi_{k}\right\|_{2}^{2}-\zeta_{Q}\delta_{Q}^{k}-\zeta_{w}\delta_{w}^{k}$ (53)
@@
$I_{w}=$ $\ \frac{3\beta\left|\mathcal{S}\right|C_{rx}^{2}}{2}+\zeta_{w}\left(-\frac{\eta\left(1-\gamma\right)^{2}}{2}+4\eta^{2}\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}\right)$
$I_{Q}=$ $\ 2\gamma^{2N}\left(\zeta_{Q}+\frac{4\eta^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\zeta_{w}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)-\zeta_{Q}$
@@
$I_{\widehat{\phi}}=$ $\ \frac{L_{\phi}^{M}}{2}\beta^{2}+2\beta^{2}L_{w}^{2}\zeta_{w}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)-\frac{\beta}{2}$
$+2\gamma^{2N}\left(\zeta_{Q}+\frac{4\eta^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\zeta_{w}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\beta^{2},$
@@
$\begin{cases}\zeta_{Q}=\zeta_{w}=1,\\ \frac{3\beta\left|\mathcal{S}\right|C_{rx}^{2}}{2\eta\left(1-\gamma\right)^{2}}<\frac{1}{2}-4\eta\left|\mathcal{S}\right|^{2}\frac{\left(1+\gamma\right)^{2}}{\left(1-\gamma\right)^{2}},\\ \beta<\frac{\tau^{2}}{8}\left(L_{\widehat{\phi}\pi}^{2}+\frac{\left|\mathcal{S}\right|^{2}\left|\mathcal{A}\right|^{3}\left(1+\gamma\right)C_{f\pi}^{2}C_{rx}^{2}}{\left(1-\gamma\right)^{2}}\right)^{-1},\\ \gamma^{2N}<\frac{1}{8}\left(1+\frac{4\eta^{2}C_{\sigma\pi}^{2}}{\tau^{2}}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)^{-1},\\ \frac{L_{\phi}^{M}}{2}\beta+2\beta L_{w}^{2}\left(1+\frac{1}{\eta\left(1-\gamma\right)^{2}}\right)+\frac{1}{4}\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\beta<\frac{1}{2}.\end{cases}$ (55)
Set the step size $\beta=\rho\eta$ with $\rho>0$. More precisely, we present the parameter configuration to guarantee (55).
$\begin{cases}\zeta_{Q}=\zeta_{w}=1,\\ \eta<\min\left\{1,\frac{\left(1-\gamma\right)^{2}}{16\left|\mathcal{S}\right|^{2}\left(1+\gamma\right)^{2}}\right\},\\ \rho<\min\left\{\frac{\left(1-\gamma\right)^{2}}{6\left|\mathcal{S}\right|C_{rx}^{2}},\frac{\left(1-\gamma\right)^{2}}{8L_{w}^{2}}\right\},\\ \beta<\min\left\{\frac{\tau^{2}}{8}\left(L_{\widehat{\phi}\pi}^{2}+\frac{\left|\mathcal{S}\right|^{2}\left|\mathcal{A}\right|^{3}\left(1+\gamma\right)C_{f\pi}^{2}C_{rx}^{2}}{\left(1-\gamma\right)^{2}}\right)^{-1},\frac{1}{4}\left(\frac{L_{\phi}^{M}}{2}+2L_{w}^{2}+\frac{1}{4}\left(\frac{C_{rx}}{1-\gamma}\right)^{2}\right)^{-1}\right\},\\ \gamma^{2N}<\frac{1}{8}\left(1+\frac{4C_{\sigma\pi}^{2}}{\tau^{2}}\left(\frac{1}{\left(1-\gamma\right)^{2}}+4\right)\right)^{-1},\end{cases}$ (56)
@@ Appendix H ANALYSIS OF SOBIRL
The next proposition, generalized from Lemma 1 in (Chakraborty et al.,, 2024), considers the distribution $\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi\right)$, with each trajectory $d_{i}=\{\left({s_{h}^{i},a_{h}^{i}}\right)\}_{h=0}^{H-1}$ ($i=1,2,\ldots,I$) sampled from the trajectory distribution $\rho\left({d;\pi}\right)$, i.e.,
@@
Denote the total variation between distributions by $D_{F}\left({\cdot,\cdot}\right)$. For any trajectory tuple $\left(d_{1},d_{2},\ldots,d_{I}\right)$, with each trajectory holding a finite horizon $H$, we have
$D_{F}\left(\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi^{1}\right),\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi^{2}\right)\right)\leq\frac{HI}{2}\sqrt{\left|\mathcal{A}\right|}\left\|\pi^{1}-\pi^{2}\right\|_{2}.$ (57)
@@
$D_{F}\left(\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi^{*}(x_{1})\right),\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi^{*}(x_{2})\right)\right)\leq\frac{HIL_{\pi}}{2}\sqrt{\left|\mathcal{A}\right|}\left\|x_{1}-x_{2}\right\|_{2}.$
@@
Note that $\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi\right)$ is the product measure of $\rho\left(d_{i};\pi\right),\,\left(i=1,2,\ldots,I\right)$, i.e., $\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi\right)=\rho\left(d_{1};\pi\right)\times\rho\left(d_{2};\pi\right)\times\cdots\times\rho\left(d_{I};\pi\right)$. By the total variation inequality for the product measure,
$D_{F}\left(\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi^{1}\right),\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi^{2}\right)\right)\leq$ $\sum_{i=1}^{I}{D_{F}\left(\rho\left(d_{i};\pi^{1}\right),\rho\left(d_{i};\pi^{2}\right)\right)}$
@@
If the vector-valued function $z\left({d_{1},d_{2},\ldots,d_{I};x}\right)\in\mathbb{R}^{n}$ is bounded by $C$, i.e., $\left\|{z\left({d_{1},d_{2},\ldots,d_{I};x}\right)}\right\|_{2}\leq C$, then for any trajectory tuple $\left(d_{1},d_{2},\ldots,d_{I}\right)$, with each trajectory holding a finite horizon $H$, and policies $\pi_{1},\pi_{2}$, it holds
@@
We resort to the property of the Integral Probability Metric (IPM) (Sriperumbudur et al.,, 2009),
@@
$\left|b^{\top}\left(\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{1}\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{2}\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]\right)\right|$
$\leq$ $\ 2C\left\|b\right\|_{2}D_{F}\left(\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi_{1}\right),\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi_{2}\right)\right)$
@@
$\left\|{\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{1}\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi_{2}\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]}\right\|_{2}$
$\leq$ $\ 2C\times D_{F}\left(\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi_{1}\right),\rho\left(\left(d_{1},d_{2},\ldots,d_{I}\right);\pi_{2}\right)\right).$ (64)
@@
$Z\left(x_{1}\right)-Z\left(x_{2}\right)=$ $\left(\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{1}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]\right)$
$+\left(\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{2}\right)\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]\right).$
@@
$\left\|\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{1}\right)\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]-\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x_{2}\right)\right)}\left[z\left(d_{1},d_{2},\ldots,d_{I};x_{2}\right)\right]\right\|\leq CHIL_{\pi}\sqrt{\left|{\mathcal{A}}\right|}\left\|x_{1}-x_{2}\right\|_{2}.$
@@
The following proposition measures the quality of the implicit differentiation estimators. A similar result is introduced as an assumption in (Shen et al.,, 2024) (Lemma 18, condition (c)). We claim that it is satisfied in our setting.
@@
Under Assumption 7.3, for any upper-level variable $x\in\mathbb{R}^{n}$ and policies $\pi^{1},\pi^{2}$, we have
$\max_{s,a}\left\|\widetilde{\nabla}Q_{sa}\left({x,\pi^{1}}\right)-\widetilde{\nabla}Q_{sa}\left(x,\pi^{2}\right)\right\|_{2}$ $\leq\frac{(1+\gamma)C_{rx}\sqrt{\left|\mathcal{A}\right|}}{\left(1-\gamma\right)^{2}}\left\|\pi^{1}-\pi^{2}\right\|_{2},$
@@
Under the Assumptions 7.2 and 7.3, $\nabla\phi(x)$ is $L_{\phi}$-Lipschitz continuous, i.e., for any $x_{1},x_{2}\in\mathbb{R}^{n}$,
@@
$\nabla\phi(x)=$ $\ \mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[\nabla l\left(d_{1},d_{2},\ldots,d_{I};x\right)\right]$ (69)
$+\tau^{-1}\mathbb{E}_{d_{i}\sim\rho\left(d;\pi^{*}\left(x\right)\right)}\left[l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)\right],$
@@
$z\left({d_{1},d_{2},\ldots,d_{I};x}\right):=l\left(d_{1},d_{2},\ldots,d_{I};x\right)\left(\sum_{i}{\sum_{h}{\nabla\left(Q_{s_{h}^{i}a_{h}^{i}}^{*}\left(x\right)-V_{s_{h}^{i}}^{*}\left(x\right)\right)}}\right)$
@@
$\left\|z\left(d_{1},d_{2},\ldots,d_{I};x\right)\right\|_{2}\leq 2HIC_{l}\frac{C_{rx}}{1-\gamma},$
@@
Under Assumptions 7.2 and 7.3, for any upper-level variable $x\in\mathbb{R}^{n}$ and policies $\pi^{1},\pi^{2}$, we have